If a project's expected return is 10%, which represents a 16% return in a booming economy and a - 8% (8% loss) return in a stagnant economy, what is the probability of a stagnant economy?

A) 18.33%

B) 25.00%

C) 33.33%

D) 50.00%
Answer #1

EXPECTED RETURN = SUM(PROB X RETURN)

EXPECTED RETURN = PROB OF BOOM X RETURN + PROB OF STAGNANT X RETURN

suppose probability of boom = W, so probability of stagnant economy = 1 -W

EXPECTED RETURN = W X RETURN + (1-W) X RETURN

10% = W X (16%) + (1-W) X (-8%)

10% = W(16%) -8% + W(8%)

18% = W(24%)

W = 18/24 = 75%

1- W = 25%

ANSWER : PROB OF STAGNANT ECONOMY = 25%

ANSWER : B) 25.00% (Thumbs up please)